# Mitigating Text Toxicity with Counterfactual Generation

### v2

@@ Abstract
Toxicity mitigation consists in rephrasing text in order to remove offensive or harmful meaning. Neural natural language processing (NLP) models have been widely used to target and mitigate textual toxicity. However, existing methods fail to detoxify text while preserving the initial non-toxic meaning at the same time. In this work, we propose to apply eXplainable AI (XAI) methods to both target and mitigate textual toxicity. We propose CF-Detox to perform text detoxification by applying local feature importance, counterfactual example generation and counterfactual feature importance methods to a toxicity classifier distinguishing between toxic and non-toxic texts. We carry out text detoxification through counterfactual generation on three datasets and compare our approach to three competitors. Automatic and human evaluations show that recently developed NLP counterfactual generators can mitigate toxicity accurately while better preserving the non-toxic meaning of the initial text as compared to classical detoxification methods. Finally, we take a step back and discuss ethical considerations on how to manage the diversity of toxicity perception and the risk of malicious use of detoxification tools. This work is the first to bridge the gap between counterfactual generation and text detoxification and paves the way towards more practical applications of XAI methods.
@@ 1 Introduction
In this paper we propose to address toxicity targeting and mitigation by applying eXplainable AI (XAI) methods Molnar (2020): more precisely Local Feature Importance (LFI) Barredo Arrieta et al. (2020), counterfactual (CF) generators (see Guidotti (2022) for a survey) and Counterfactual Feature Importance (CFI) Bhan et al. (2023a). The former aims at detecting important input features to explain a model prediction. Instead, CF example generation explains a model prediction by identifying the minimal changes that enable flipping the outcome of a classifier. CFI measures token modification importance to explain label flipping related to CF explanations.
@@
We propose CF-Detox, a detoxification method based on a recently developed CF generator: TIGTEC Bhan et al. (2023b) and validated by both automatic and human experiments,
@@ 2 Background and Related Work
This section recalls the task of automatic toxicity detection and mitigation. We presents existing methods aiming to detoxify text using neural NLP models. We then the XAI principles used in the next section to perform toxicity detection and mitigation. In the following, we employ text detoxification and toxicity mitigation interchangeably.
@@ 2.1 Automatic Toxicity Processing background
2.1.1 Definition and objective
Textual toxicity can be defined in multiple ways Fortuna and Nunes (2018) and can take various forms, such as rude, offensive or hateful speech, potentially causing online harm to isolated people or minority groups Thomas et al. (2021). Automatic toxicity processing can essentially take two forms: detection and mitigation. Toxicity detection can be either based on prior knowledge (vocabulary, regex) or obtained from a fine tuned toxicity classifier mapping an input text representation space to an output space to distinguish toxic and non-toxic texts Cook et al. (2023). Training language models to perform this classification is difficult, as it requires access to datasets labeled based on an idiosyncratic definition of toxicity derived from human annotators Jigsaw (2018). Toxicity mitigation consists in rewriting a toxic text while preserving the non-toxic meaning. This task is even more difficult because it requires disentangling toxic and non-toxic meanings to modify the former plausibly while preserving the latter.
2.1.2 Evaluation criteria of text detoxification
@@ 2.2 Toxicity Mitigation with Neural NLP Models
2.2.1 Text Style Transfer
@@ 2.3 XAI for NLP
2.3.1 Local Feature Importance
A Local Feature Importance (LFI) function explains a prediction by a vector where is the contribution of the th token to the prediction. The higher the contribution, the more important the token is to explain the prediction of the classifier . Three types of LFI methods can be distinguished: perturbation-based such as KernelSHAP Lundberg and Lee (2017), gradient-based such as Integrated gradients Sundararajan et al. (2017) and attention-based such as self-attention in case of a Transformer classifier Bhan et al. (2023c).
2.3.2 Counterfactual explanations
@@
2.3.3 Counterfactual Feature Importance
A Counterfactual Feature Importance (CFI) operator computes token substitution importance between an instance of interest and its related CF explanation Bhan et al. (2023a). In the same way as common LFI, is the contribution of the -th token modification. CFI highlights the most important token modifications, making non-sparse CF explanations easier to read and more intelligible. A simple but effective way to compute CFI is to apply Integrated gradients to explain the difference between and , by setting the baseline to the initial instance of interest .
@@ 3.2 Epistemic Similarities Between Text Detoxification and Counterfactual Generation
Thirdly, methods detoxifying text and generating CFs can be grouped into two similar categories. As mentioned in Bhan et al. (2023b), textual CF generators can be of two types: Text editing heuristics and Counterfactual generation with large language models.
Text editing heuristics is a family of CF generation methods that builds CF examples by slightly modifying the input text whose prediction is to be explained. Important tokens are targeted and modified with mask language models to switch the outcome of the classifier, making this approach very similar to the M&R way of detoxifying text. Text editing heuristics methods differ in the way they target important tokens and the language model used to modify the initial text. Regarding the former methods, they mostly target tokens to be modified by applying LFI methods to the classifier that has to be explained. For example, CLOSS Fern and Pope (2021) applies Ablation, MiCE Ross et al. (2021) gradient-based approaches, CREST Treviso et al. (2023) leverages rationalization methods, and TIGTEC Bhan et al. (2023b) employs KernelSHAP or Self-attention. Next, mask language models used to perform mask infilling are essentially T5, RoBERTa or BERT. This way, M&R methods and Text editing heuristics differ only in the way they target toxic POS. As mentioned in the previous section, the former detect tokens to change either by the use of prior knowledge (vocabulary), intrinsically interpretable model (logistic model) or expert-anti-expert disagreement, whereas Text editing heuristics apply LFI methods to a toxicity classifier. Finally, text is modified using the same kind of NLP encoder. We propose to group these two kinds of approaches together under the name of "target-then-replace" methods. Figure 2 shows the raw target-then-replace CF detoxification process, where the detoxification consists in performing the two following token changes: f**k hell and you they, leading to a text certainly detoxified, but not perfectly content preserving.
@@ 4.1.1 Datasets
4.1.2 Counterfactual generator and competitors
We instantiate our method proposed in the previous section by setting as CF generation method TIGTEC Bhan et al. (2023b). The rationale behind this choice is that TIGTEC achieves the best compromise in terms of accuracy, content preservation and plausibility when compared with other major competitors Bhan et al. (2023b). TIGTEC is a target-then-replace textual CF generator that implements several LFI methods to target important tokens to be changed. It iteratively masks and replaces tokens with a BERT mask language model following a tree search policy based on beam search to minimize a cost function. We use the following settings to run TIGTEC, we first train a neural classifier (BERT or DistilBERT) on a toxic task dataset from Kaggle222https://www.kaggle.com/datasets/rounak02/imported-data to learn to distinguish toxic texts from non-toxic ones. The classifier performance after training is respectively 94.1% and 90.2% for BERT and DistilBERT. This way, CF text detoxification is performed with classifiers that have been trained on a different dataset from the ones used for the evaluation.
Toxicity mitigation is then performed by generating CF examples starting from toxic texts in order to reach a non-toxic state. In the following, we call CF-Detox our new toxicity mitigation method based on CF generation. The cost function driving CF search is defined such as: with where is the probability score of being non-toxic with respect to the trained classifier, is a cosine similarity measure between text embeddings obtained from Sentence Transformers and is a hyperparameter set to 0.3. For each neural model, CF-Detox is run in three different versions, targeting toxic POS with either KernelSHAP (kshap), Self-attention (attention) or Integrated gradients (IG). Self-attention is aggregated as in the original TIGTEC paper by averaging all the attention coefficients related to the CLS token over the attention heads in the last layer of the classifier.
4.1.3 Automatic evaluation
4.1.4 Human-grounded evaluation
In addition to automatic evaluation, we perform a human-grounded experiment to compare CF-Detox to MaRCo, CondBERT and ParaGeDi in terms of toxicity mitigation performance. It consists in asking 5 annotators to rank detoxified texts by toxicity level obtained by applying CF-Detox, MaRCo, CondBERT and ParaGeDi on 20 randomly selected texts from each dataset. The order of appearance of the toxicity mitigation methods and the dataset is randomized, so that there is no spatial bias in information processing. Annotators can rank texts at the same level if necessary.
Before running the experiment, annotators are given the same instructions. To make sure that they annotate based on the same common knowledge, we define the textual toxicity as "violent, aggressive or offensive language that may focus on a specific person or group of people sharing a common property. This common property can be gender, sexual orientation, ethnicity, age, religion or political affiliation.". Annotators all have a MSc degree in data analytics or machine learning and have a good knowledge of English.
4.2 Results
4.2.1 Global results
@@
4.2.2 Human evaluation
This way, automatic and human evaluation underpin the fact that CF-Detox offers another possible compromise between toxicity, meaning preservation and text plausibility as compared to other state-of-the-art existing methods.
4.2.3 Ablation study
@@ 4.2.4 CFI enhancement
We refine detoxified texts using CFI following the target-then-replace-then-refine methodology previously introduced. Table 3 shows similarity and sparsity growth rate obtained from CFs obtained by running CF-Detox with KernelSHAP to target toxic POS. On average, CFI increases both similarity and sparsity across all datasets, when CF-Detox is either based on BERT or DistilBERT. However, the gains are higher with the DynaHate dataset (respectively +4.7% and +4.8% uplift in sparsity for BERT and DistilBERT) containing the most toxic texts as compared to MAgr and SBF.
@@ 7.4 Personal Data
Each voluntary participant to the human evaluation signed an informed consent form outlining the project’s purpose and details. The data were anonymized and processed only by the authors and stored in accordance with the General Data Protection Regulation (GDPR). It was possible to stop at any time. The Consent form used is anonymized and presented in the supplementary material. This study was conducted for research purposes only.
7.5 Ecological impact
The use of NLP neural models to mitigate toxicity induces a non-negligible carbon consumption Crawford (2021). Most of the detoxification methods rely on neural NLP models for text generation under control. For instance, both MarCo and TIGTEC use three language models to perform toxicity mitigation. These language models have several million (or billion) parameters, which often require the use of high-emission clusters to process large quantities of data. These methods have to be used with caution and alternatives to neural NLP have to be considered whenever possible.
8 Discussion
In this work we showed that XAI methods can be applied to a toxicity classifier to target toxic POS with LFI and mitigate toxicity with counterfactual generation. Counterfactual detoxification with the TIGTEC counterfactual generator enables us to find a new compromise in terms of toxicity lowering, content preservation, and textual plausibility.
Counterfactual toxicity mitigation is highly dependent on the toxicity classifier used to steer detoxification. If is unreliable and only performs well on its training set, the risk of incorrectly indicating that the text has been detoxified is high. Therefore, the choice of and the toxicity training data must be made with caution to avoid incorrect toxicity assessments during detoxification. The classifier can be fine-tuned on a more specific dataset if the detoxification task is related to a precise kind of toxicity, such as racism or sexism.
Counterfactual detoxification has been tested by applying three different versions of TIGTEC with SHAP, Integrated gradients, or Self-attention. There is a wide range of other LFI methods that could be used to target important tokens to explain a toxicity prediction, such as LIME Ribeiro et al. (2016) or DeepLift Shrikumar et al. (2017). Besides, other counterfactual generators such as MiCE Ross et al. (2021), CREST Treviso et al. (2023) or CLOSS Fern and Pope (2021) could be used to perform toxicity mitigation. We believe that these counterfactual generation methods could lead to other levels of compromise between toxicity lowering, text plausibility and content preservation.
This paper is the first to show the extent to which fields such as automatic toxicity processing and explainable AI, which have developed in parallel, actually share many similarities and can be mutually beneficial.
@@ 9 Conclusion
This paper formalized how LFI and counterfactual generation methods can be used to target textual toxic content and perform toxicity mitigation. CF-Detox leads to competitive results, with state-of-the-art performance in terms of content preservation while accurately detoxifying text and generating plausible text. CF-Detox is versatile since it can be used with various types of LFI methods (such as attention, gradient, and perturbation) to target toxic POS. This work is the first attempt to recognize the systemic similarity of these tasks and address detoxification through counterfactual generation. While counterfactual toxicity mitigation may yield competitive results, it also poses risks in terms of malicious use. In particular, counterfactual detoxifiers could be hijacked to generate adversarial attacks or toxic content from non toxic sources, which calls for the implementation of robust good practices.
Appendix A Appendix
A.1 Human-grounded protocol participant consent form
Each participant signed an informed consent form containing the project purpose and details and the intended use of the data they would generate. Figure 6 shows an anonymized version of the consent form.

### v3

@@ Abstract
Toxicity mitigation consists in rephrasing text in order to remove offensive or harmful meaning. Neural natural language processing (NLP) models have been widely used to target and mitigate textual toxicity. However, existing methods fail to detoxify text while preserving the initial non-toxic meaning at the same time. In this work, we propose to apply eXplainable AI (XAI) methods to both target and mitigate textual toxicity. We propose CF-Detox to perform text detoxification by applying local feature importance, counterfactual example generation and counterfactual feature importance methods to a toxicity classifier distinguishing between toxic and non-toxic texts. We carry out text detoxification through counterfactual generation on three datasets and compare our approach to three competitors. Automatic and human evaluations show that recently developed NLP counterfactual generators lead to competitive results in toxicity mitigation. This work is the first to bridge the gap between counterfactual generation and text detoxification and paves the way towards more practical applications of XAI methods.
@@ 1 Introduction
In this paper we propose to address toxicity targeting and mitigation by applying eXplainable AI (XAI) methods Molnar (2020): more precisely Local Feature Importance (LFI) Barredo Arrieta et al. (2020), counterfactual (CF) generators (see Guidotti (2022) for a survey) and Counterfactual Feature Importance (CFI) Bhan et al. (2023b). The former aims at detecting important input features to explain a model prediction. Instead, CF example generation explains a model prediction by identifying the minimal changes that enable flipping the outcome of a classifier. CFI measures token modification importance to explain label flipping related to CF explanations.
@@
2.
We propose CF-Detox, a detoxification method based on a recently developed CF generator: TIGTEC Bhan et al. (2023c) and validated by both automatic and human experiments,
3.
@@ 2 Background and Related Work
This section recalls the task of automatic toxicity detection and mitigation. We presents existing methods aiming to detoxify text using neural NLP models. We then introduce the XAI principles used in the next section to perform toxicity detection and mitigation. In the following, we employ text detoxification and toxicity mitigation interchangeably.
@@ 2.1 Automatic Toxicity Processing background
2.1.1 Definition and objective.
Textual toxicity can be defined in multiple ways Fortuna and Nunes (2018) and can take various forms, such as rude, offensive or hateful speech, potentially causing online harm to isolated people, minority groups or different ethnic, religious or racial groups Thomas et al. (2021). Automatic toxicity processing can essentially take two forms: detection and mitigation. Toxicity detection can be either based on prior knowledge (vocabulary, regex) or obtained from a fine tuned toxicity classifier mapping an input text representation space to an output space to distinguish toxic and non-toxic texts Cook et al. (2023). Training language models to perform this classification is difficult, as it requires access to datasets labeled based on an idiosyncratic definition of toxicity derived from human annotators Jigsaw (2018). Toxicity mitigation consists in rewriting a toxic text while preserving the non-toxic meaning. This task is even more difficult because it requires disentangling toxic and non-toxic meanings to modify the former plausibly while preserving the latter.
2.1.2 Evaluation criteria of text detoxification.
@@ 2.2 Toxicity Mitigation with Neural NLP Models
2.2.1 Text Style Transfer.
@@ 2.3 XAI for NLP
2.3.1 Local Feature Importance.
A Local Feature Importance (LFI) function explains a prediction by a vector where is the contribution of the th token to the prediction. The higher the contribution, the more important the token is to explain the prediction of the classifier . Three types of LFI methods can be distinguished: perturbation-based such as KernelSHAP Lundberg and Lee (2017), gradient-based such as Integrated gradients Sundararajan et al. (2017) and attention-based such as self-attention in case of a Transformer classifier Bhan et al. (2023a).
2.3.2 Counterfactual explanations.
@@
2.3.3 Counterfactual Feature Importance.
A Counterfactual Feature Importance (CFI) operator computes token substitution importance between an instance of interest and its related CF explanation Bhan et al. (2023b). In the same way as common LFI, is the contribution of the -th token modification. CFI highlights the most important token modifications, making non-sparse CF explanations easier to read and more intelligible. A simple but effective way to compute CFI is to apply Integrated gradients to explain the difference between and , by setting the baseline to the initial instance of interest .
@@ 3.2 Epistemic Similarities Between Text Detoxification and Counterfactual Generation
Thirdly, methods detoxifying text and generating CFs can be grouped into two similar categories. As mentioned in Bhan et al. (2023c), textual CF generators can be of two types: Text editing heuristics and Counterfactual generation with large language models.
Text editing heuristics is a family of CF generation methods that builds CF examples by slightly modifying the input text whose prediction is to be explained. Important tokens are targeted and modified with mask language models to switch the outcome of the classifier, making this approach very similar to the M&R way of detoxifying text. Text editing heuristics methods differ in the way they target important tokens and the language model used to modify the initial text. Regarding the former methods, they mostly target tokens to be modified by applying LFI methods to the classifier that has to be explained. For example, CLOSS Fern and Pope (2021) applies Ablation, MiCE Ross et al. (2021) gradient-based approaches, CREST Treviso et al. (2023) leverages rationalization methods, and TIGTEC Bhan et al. (2023c) employs KernelSHAP or Self-attention. Next, mask language models used to perform mask infilling are essentially T5, RoBERTa or BERT. This way, M&R methods and Text editing heuristics differ only in the way they target toxic POS. As mentioned in the previous section, the former detect tokens to change either by the use of prior knowledge (vocabulary), intrinsically interpretable model (logistic model) or expert-anti-expert disagreement, whereas Text editing heuristics apply LFI methods to a toxicity classifier. Finally, text is modified using the same kind of NLP encoder. We propose to group these two kinds of approaches together under the name of "target-then-replace" methods. Figure 2 shows the raw target-then-replace CF detoxification process, where the detoxification consists in performing the two following token changes: f**k heck and you they, leading to a text certainly detoxified, but not perfectly content preserving.
@@ 4.1 Experimental Protocol
4.1.2 Counterfactual generator and competitors.
We instantiate our method proposed in the previous section by setting as CF generation method TIGTEC Bhan et al. (2023c). The rationale behind this choice is that TIGTEC achieves the best compromise in terms of accuracy, content preservation and plausibility when compared with other major competitors Bhan et al. (2023c). TIGTEC is a target-then-replace textual CF generator that implements several LFI methods to target important tokens to be changed. It iteratively masks and replaces tokens with a BERT mask language model following a tree search policy based on beam search to minimize a cost function. We use the following settings to run TIGTEC, we first train a neural classifier (BERT or DistilBERT) on a toxic task dataset from Kaggle222https://www.kaggle.com/datasets/rounak02/imported-data to learn to distinguish toxic texts from non-toxic ones. The classifier performance after training is respectively 94.1% and 90.2% for BERT and DistilBERT. This way, CF text detoxification is performed with classifiers that have been trained on a different dataset from the ones used for the evaluation.
Toxicity mitigation is then performed by generating CF examples starting from toxic texts in order to reach a non-toxic state. In the following, we call CF-Detox our new toxicity mitigation method based on CF generation. The cost function driving CF search is defined as in the TIGTEC paper as with where is the probability score of being non-toxic with respect to the trained classifier, is a cosine similarity measure between text embeddings obtained from Sentence Transformers and is a hyperparameter set to 0.3. This way, minimizing the cost function implies reaching a counterfactual state while staying semanticly close to the initial instance. The TIGTEC counterfactual search is performed in a beam search fashion, with . For each neural model, CF-Detox is run in three different versions, targeting toxic POS with either KernelSHAP (kshap), Self-attention (attention) or Integrated gradients (IG). Self-attention is aggregated as in the original TIGTEC paper by averaging all the attention coefficients related to the CLS token over the attention heads in the last layer of the classifier.
4.1.3 Automatic evaluation.
4.1.4 Human-grounded evaluation.
In addition to automatic evaluation, we conduct a human-grounded experiment to compare CF-Detox, MaRCo, CondBERT and ParaGeDi in terms of toxicity mitigation performance. It consists in asking 5 annotators to rank detoxified texts by toxicity level obtained by applying CF-Detox, MaRCo, CondBERT and ParaGeDi on 20 randomly selected texts from each dataset. The order of appearance of the toxicity mitigation methods and the dataset is randomized, so that there is no spatial bias in information processing. Annotators can rank texts at the same level if necessary.
Before running the experiment, annotators are given the same instructions. To make sure that they annotate based on the same common knowledge, we define the textual toxicity as "violent, aggressive or offensive language that may focus on a specific person or group of people sharing a common property. This common property can be gender, sexual orientation, ethnicity, age, religion or political affiliation.". Annotators all have a MSc degree in data analytics or machine learning and have a good knowledge of English. None of the annotators are authors.
4.2 Results
4.2.1 Global results.
@@ 4.2 Results
4.2.2 Human evaluation.
This way, automatic and human evaluation underpin that CF-Detox offers another possible compromise between toxicity, meaning preservation and text plausibility as compared to other state-of-the-art existing methods.
@@
4.2.3 Ablation study.
@@
We refine detoxified texts using CFI following the target-then-replace-then-refine methodology previously introduced in Section 3.3. Table 3 shows similarity and sparsity growth rate obtained from CFs obtained by running CF-Detox with KernelSHAP to target toxic POS. On average, CFI increases both similarity and sparsity across all datasets, when CF-Detox is either based on BERT or DistilBERT. However, the gains are higher with the DynaHate dataset (respectively +4.7% and +4.8% uplift in sparsity for BERT and DistilBERT) containing the most toxic texts as compared to MAgr and SBF.
@@ 6.4 Personal Data
Each voluntary participant to the human evaluation signed an informed consent form outlining the project’s purpose and details. The data were anonymized and processed only by the authors and stored in accordance with the General Data Protection Regulation (GDPR). It was possible to stop at any time. The Consent form used is anonymized and presented in Appendix 0.A. This study was conducted for research purposes only.
@@ 7 Conclusion
Appendix 0.A Appendix
0.A.1 Consent form
0.A.2 Implementation Details
The code used to run the experiments is available on the github related to the TIGTEC paper333https://github.com/milanbhan/tigtec.
Pretrained language models
The library used to import pretrained language model to be finetuned to distinguish toxic and non-toxic content is transformer. In particular, the backbone version of BERT is bert-base-uncased the one of DistilBERT is distilbert-base-uncased. The library used to import the Sentence Transformer is sentence_transformers and the model backbone is paraphrase-MiniLM-L6-v2. The backbone of the pre-trained toxicity classifier is toxic-bert.